OpenStudy (anonymous):

Multiply (x − 4)(x2 − 3x + 5) x3 + 5x2 + 5x − 20 x3 + 2x2 + 8x − 20 x3 − x2 + 7x − 20 x3 − 7x2 + 17x − 20

OpenStudy (abdela25):

(x − 4)(x2 − 3x + 5) X = 4 x3 + 5x2 + 5x − 20 X = 1.418 assuming x3 means X^3 x3 − x2 + 7x − 20 X = 2.12 also assuming X3 means X^3 x3 − 7x2 + 17x − 20 X = 4 assuming X3 menas X^3 do you know how i did this?

OpenStudy (anonymous):

np

OpenStudy (anonymous):

I can give u metal if i don't know the answer

OpenStudy (anonymous):

I can not

OpenStudy (anonymous):

Hi. So do this, we need to foil. Let's separate both equations: (x) * (x^2-3x+5) Now distribute the x. Do you know how to do that?

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

so will be x^2-3x^2+5x

OpenStudy (anonymous):

Almost, remember that x*x^2=x^3, but everything else is right. :) Now the second one: -4*(x^2-3x+5)=?

OpenStudy (anonymous):

4x^2+12x-20

OpenStudy (anonymous):

so the final answers is x^2-7x+17x-20

OpenStudy (anonymous):

Now remember the - in -4x^2 x^3-3x^2-4x^2+5x+12x-20 The final answer is x^3-7x^2+17x-20 You got the answer right, just don't forget the ^signs :)

OpenStudy (anonymous):

thank you help me on this one last one

OpenStudy (anonymous):

Simplify (8j3 + 5j2 − 3) − (5j3 + 7j2 − 12j + 7) 3j3 + 12j2 − 15j + 10 3j3 + 12j2 + 12j + 4 3j3 − 2j2 + 12j − 10 3j3 − 2j2 − 12j + 4

OpenStudy (anonymous):

so first we need to distribute the negative: -(5j3+7j2-12j+7)=?

OpenStudy (anonymous):

change the sing

OpenStudy (anonymous):

Yes, when you distribute a negative, you basically change the sign to the opposite sign

OpenStudy (anonymous):

-5j^3-7j^2+12j-7

OpenStudy (anonymous):

Yes, now put the whole question together: 8j3+5j2-3-5j3-7j2+12j-7 Now combine like terms. What will be the final answer?

OpenStudy (anonymous):

3j^3-2j^2-15-7

OpenStudy (anonymous):

3j3 is right, -2j2 is right, but if you notice in the problem, there is only 12j, there's no like term to combine that with, but -3 and -7 equals -10, so it will be 3j^3-2j^2+12j-10 Do you understand it?

OpenStudy (anonymous):

thanks what about this one Rationalize the denominator of square root of negative 16 over open parentheses 1 plus I close parentheses plus open parentheses 6 plus 3 i quantity of negative 16 plus 28 I over 33 quantity of 16 plus 28 I over 65 quantity of negative 12 plus 28 I over 40 quantity of 12 plus 28 I over 58

OpenStudy (anonymous):

Can you draw out the question so I can have a better view of how it looks?

OpenStudy (anonymous):

ok\[\sqrt{-16}/((1+i)+(6+3i)\]

OpenStudy (anonymous):

I don't know how to write it correctly

OpenStudy (anonymous):

Thank you. First thing we need to do is to simplify the denominator: 1+i+6+3i Combine like terms 1+i+6+3i=?

OpenStudy (anonymous):

7-3i^2

OpenStudy (anonymous):

It should be 7+4i because we are adding 3i and i.

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

Next, we simplify the top part. sqrt of 16 is 4, but because 16 is negative, the answer will be 4i. Now to rationalize the denominator, we have multiply the top and bottom by the conjugate of (7+4i), which is (7-4i)

OpenStudy (anonymous):

So we start with the top one: 4i(7-4I) Distribute the 4i, equals ?

OpenStudy (anonymous):

28i-16i^2

OpenStudy (anonymous):

It seen that you love math with<3

OpenStudy (anonymous):

Yes. Now the bottom: This time we foil (7+4i)(7-4i)= 49-28i+28i-16i^2 Now the simplify =? and yeah I love math lol

OpenStudy (anonymous):

so final answers is 16+28/33

OpenStudy (anonymous):

hey I an working on this assignment seen two moth and i can figurative out can you make a look at it I will really appreciated. thank you and god bless you.

OpenStudy (anonymous):

Well, the top we have 28i-16i^2, and I^2 means (-1), so -16*-1 equals 16, and the bottom is 49-16i^2 which is becomes 49+16 So let's put it all together: 28i+16/65 so the answer will be b

OpenStudy (anonymous):

i got it thanks can you check what i send you:(

OpenStudy (anonymous):

Ok so I'm looking through it, and for number 1 c, it should be x^3+3x^2+9 instead of x^2+3x^2+9

OpenStudy (anonymous):

So I'm not sure for the rest of them. Sorry :(

OpenStudy (anonymous):

ok don't worry thank for your help when i need help can I sen you a message

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@tHe_FiZiCx99 there's a link for the document that has the questions on it up.

OpenStudy (the_fizicx99):

Which question is it?

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Number 3 on the document. The one about grouping method.

OpenStudy (the_fizicx99):

Are you suppose to pick from certain factoring methods? I use the Christmas tree method to factor but when you'e doing quadratics there's a better method, don't remember the name though.

OpenStudy (anonymous):

ok here is the Hw with no numbers this is what the teachers gave me

OpenStudy (anonymous):

I do't know confuse:/

OpenStudy (anonymous):

@yulima How about I message you the answer after I ask my brother about it? He knows more about math than me :P

OpenStudy (anonymous):

ok i really appreciated.

OpenStudy (anonymous):

:)